A solid rectangular block measures 0.1m by 0.08m by 0.06m and floats freely in a liquid of density 13546kg/m3. If the depth of the liquid is 0.045m up to the block's largest side. Find the density of the block?


1
Expert's answer
2021-01-05T12:03:42-0500

"mg=F_A"

"\rho Vg=\rho_l V_l g"

"\rho V= \rho_l V_l"

"\rho h S=\rho_l h_l S"

"\rho h=\rho_l h_l"

"\rho=\rho_l \frac{h_l}{h}=13546\cdot \frac{0.045}{0.1}=6096 ~\frac{kg}{m^3}"